# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ Abstract
Power system operators and electric utility companies often charge consumers a peak demand charge when the aggregate system demand reaches its maximum, a practice known as coincident peak shaving. These charges incentivize consumers to reduce usage during critical periods, alleviating stress on electricity transmission and distribution systems, while also helping to recover the grid investment costs. In this paper, we analyze the problem through the lens of game theory, developing a model that captures how strategic consumer behavior influences overall system efficiency. Our results reveal that the coincident peak shaving game can be concave, quasi-concave with discontinuities, or non-concave with discontinuities, depending on the extent of consumers’ demand-shifting capabilities. In a two-agent, two-period framework, we derive closed-form Nash equilibrium solutions for each scenario and extend our analysis to multi-agent settings. We prove the stability of these equilibrium points and propose an algorithm to compute equilibrium outcomes across all game configurations. Furthermore, we show that while the decentralized game model achieves peak-shaving performance comparable to a centralized approach, it does so at the cost of increased anarchy. In scenarios characterized by quasi-concave and non-concave conditions, our analytical results demonstrate that anarchy grows with greater consumer flexibility and disparities in marginal shifting costs, and we explore how the number of agents affects system efficiency. Finally, numerical simulations are provided to validate our theoretical findings.
@@
Coincident peak shaving, Bottleneck game, Price of anarchy, Equilibrium stability, Demand response, Power system operation
@@ I Introduction
Managing peak electricity demand is crucial in power system transmission and distribution due to the need to maintain an instantaneous balance between supply and demand. At the transmission level, total demand must not exceed the maximum generation capacity. At the distribution level, where a tree-branched network of transformer substations steps down voltages, the load at each tier must remain below the capacity of the respective transformers. When demand exceeds the system’s generation or substation capacity, operators may be forced to disconnect portions of the grid to cut off the connecting demand to ensure the secure operation of the remaining system, potentially leading to significant financial losses and adverse impacts on the life quality [Kirschen and Strbac, 2018].
@@
With the advent of smart electric meters capable of recording time-dependent electricity usage and the widespread deployment of distributed energy resources (DER), power system operators and electric utility companies are implementing new pricing schemes to promote demand flexibility and alleviate peak demand stress. One such approach is time-of-use (ToU) tariffs [Salt River Project, 2025], in which utilities set hourly electricity prices to encourage consumers to shift consumption: increasing usage during low-priced periods (typically at night) and reducing it during high-priced periods (typically during the day). However, ToU tariffs mainly focus on recovering the cost of energy supply and shifting daily consumption, and they are less effective at curbing peak demand over longer periods, such as on a monthly basis, because the prices are not set high enough to further incentivize significant peak demand reduction [Trabish, 2016, Burger et al., 2020].
@@
Utility companies may also impose direct peak demand charges, billing consumers based on their highest power usage with a prefixed peak price over an extended period—typically a month. These charges not only motivate consumers to reduce their peak consumption but also help recover distribution grid investment and maintenance costs, as the delivered grid capacity is determined by a consumer’s peak demand [Mullendore, 2017, Shi et al., 2017]. For instance, Consolidated Edison, serving New York City, may charge more than \$40 per kW of peak demand, with the peak being determined by the highest two consecutive 15-minute intervals in any monthly billing cycle. This effectively translates to a cost of approximately \$80 per kWh during peak periods, which is significantly higher than the standard energy supply rate of around 20 to 30 cents per kWh [Consolidated Edison, 2025a].
@@
Yet, allocating peak demand charges becomes non-trivial when multiple consumers share a capacity bottleneck, such as a constrained substation, because the system’s peak does not directly correspond to each individual consumer’s peak [Chen et al., 2020b]. Charging a consumer for their peak demand may be inequitable if the overall system is not under peak stress, and such charges might not accurately represent the true system peak demand. This issue is particularly critical for utility companies, which are often publicly owned or regulated private monopolies with profit margins capped by their investment costs. Therefore, the revenue from peak demand charges should accurately reflect grid investment expenses based on the actual delivery capacity.
@@
Coincident peak (CP) demand charges bill consumers based on their usage during the system’s peak period, offering an alternative that more accurately reflects the system condition and the corresponding investment costs. These charges are added to future electricity bills; for example, Texas’s 4CP mechanism applies charges based on a consumer’s average demand during the Electric Reliability Council of Texas (ERCOT) system’s peak hour in each of the four summer months (from June to September), with these charges affecting the following year’s bills. Notably, 4CP charges can represent up to 30% of an organization’s monthly electricity costs [Ögelman, 2016]. In 2023, the 4CP charge was approximately \$67/kWh [Grid Status, 2024], with the total coincident peak demand across the ERCOT region reaching 83.6 GW [ERCOT, 2025b], translating into an estimated \$5.6 billion payments. According to ERCOT’s annual demand response report, the 4CP mechanism facilitated a 3,500 MW demand reduction during peak periods, resulting in savings of approximately \$234.5 million, or 4.2% of the total 4CP-related costs [Potomac Economics, 2023]. A significant challenge, however, is that the peak period is identified only in hindsight. While numerous academic and industrial efforts have focused on accurately predicting the system’s peak period [Patrick and Wolak, 2001, Dowling et al., 2018, CPower, 2016], these studies often overlook the interactions among consumers, whose collective behavior determines the system peak. This gap naturally motivates a game-based framework and raises a critical question: Is a game-based framework workable for the CP shaving problem, and how does strategic consumer behavior affect system efficiency? Addressing this question is the main focus of this paper.
@@ I-A Summary of main contributions and implications
In this work, we formulate a novel game framework for the CP shaving problem. As a first step toward applying a game-theoretic model to address CP shaving challenges, we consider a simplified two-agent, two-period setting, with extensions to more than two agents discussed in Section VI. The two agents represent clusters of customers or dominant oligopoly customers. The rationale for the two-period setting stems from its ability to represent a peak period and an off-peak period [Anunrojwong et al., 2024], as identified from historical market data, including demand shifting patterns and price signals. For example, power system peak demand typically occurs in the early evening, with off-peak periods covering the rest of the day. In the 4CP program in Texas, historical records in 2024 suggest that certain periods, such as around 17:00, are more likely to experience CP events [ERCOT, 2025b]. Our model focuses on minimizing system peak demand while deemphasizing off-peak performance, aligning with the primary objective of operators and utilities. This simplified framework offers both insight and tractability for analyzing the fundamental interactions between agents and their impact on system performance. Furthermore, as agents can access the system demand profile from publicly available market information [ERCOT, 2025a], they can infer the aggregated response of other agents. We assume that each agent observes the other’s aggregated demand distribution, allowing us to isolate strategic interactions from the complexities of predicting individual valuations. By balancing simplicity with realism, this approach lays a solid foundation for understanding CP shaving strategies and agent behaviors.
@@
With our settings in place, we formulate the CP game model with a fixed CP charge price and an individual penalty term for demand shifting to determine the cost, where agents compete to determine the CP period and their corresponding demand strategies. The structure of the CP game model is influenced by the agents’ demand-shifting capabilities (flexibility), which are determined by their shifting costs and baseline demands. Agents with larger baseline demand differences between two periods and lower shifting costs exhibit greater flexibility. Depending on these factors, the model may exhibit concave, quasi-concave with discontinuities, or non-concave with discontinuities characteristics. By contrast, the conventional centralized peak shaving model is formulated as a straightforward convex optimization problem.
@@
We then analytically derive the pure strategy Nash equilibrium (NE) for all CP game structures under the two-agent, two-period setting. Our analysis reveals that the NE is determined by the system’s unbalanced demand and the agents’ shifting capabilities. By treating the system as a switched dynamic system, we prove that the equilibrium of the game model is globally uniformly asymptotically stable, regardless of switching, provided all agents’ demand shifting is within the baseline limit. Furthermore, we demonstrate that a gradient-based algorithm with an update rule serving as a finite difference approximation of the asymptotically stable process can compute the equilibrium point.
@@
We analyze the impact of gaming agents’ strategic behaviors by examining the peak shaving effectiveness and the price of anarchy (PoA). Our findings analytically indicate that the peak shaving effectiveness of the game model matches that of the centralized model but comes at the cost of economic inefficiency (anarchy), except for the concave game, where the outcomes are fully equivalent. For quasiconcave and non-concave games, we prove that the PoA increases with the disparities among agents, as measured by their marginal shifting costs. Additionally, under identical system conditions, we show that greater agent flexibility exacerbates PoA.
@@
We also generalized our findings with more than two agents, we prove the unique NE under concave and quasiconcave conditions. For non-concave CP games, we show that system demand can still be balanced across two periods and derive the NE at the group level of agents. We also establish the stability of the equilibrium point and validate the effectiveness of the gradient-based algorithm. The NE reveals equivalent peak shaving effectiveness between the game and the centralized model while showing that changes in the number of agents influence the game type and, consequently, the PoA.
For industry practitioners and policymakers, our paper offers the following takeaways
@@
  • We show that the game model with a stable NE is computationally efficient, and applying this model to the CP shaving problem does not compromise peak shaving performance. Thus, operators can adopt the game model without concerns about its impact on CP shaving effectiveness.
@@ I-B Related works
To precisely position our study within the gap in the existing literature, we elaborate on our work from three perspectives closely related to our CP shaving problem: (i) Load management, (ii) CP demand charge, and (iii) Related game formulations.
@@
Load management. Load management is a border concept covering peak demand reduction, load shifting, and demand response. Many studies, from the utility’s perspective, formulate optimization problems to design pricing mechanisms - such as tariffs or peak demand charges - aimed at minimizing convex costs or maximizing concave profits [Siano, 2014]. These pricing design approaches often rely on consumers’ price response behavior, typically leading to a two-layer framework (or Stackelberg game model): the upper layer determines the pricing strategy, while the lower layer models demand based on consumer response behavior [Chen et al., 2020a]. The lower layer may involve constructing utility functions to represent consumer preferences [Pennings and Smidts, 2003, Chen and Xu, 2024] or applying data-driven methods to learn behavior from historical price and consumption data [Jain et al., 2017]. From a demand-side perspective, the key challenge lies in distinguishing between baseline demand and observed demand to better understand consumers’ demand reduction decision-making behavior [Agrawal and Yücel, 2022]. Beyond individual consumers, the interactive influence of demand strategies has been widely studied through non-cooperative game formulations and generally achieves peak shaving [Joshi et al., 2021]. However, while these load management approaches contribute to peak demand reduction, they primarily focus on maximizing the profit of different market participants rather than directly addressing CP demand reduction in accordance with CP charging mechanisms.
@@
CP demand charge. Academic studies and industrial solutions for CP shaving generally focus on predicting the timing of the system peak and making decisions from the consumers’ perspective. Prediction involves estimating the CP period as a probabilistic distribution, often using machine learning tools that leverage input features such as historical demand and weather conditions [Carmona et al., 2024, Dowling et al., 2018]. Following this, decisions are made based on the prediction results. Industrial solutions primarily emphasize short-term prediction, employing auto-regressive methods to iteratively update models and issue warning signals to consumers [CPower, 2016]. Academic solutions typically frame the problem as a scheduling task, making deliberate decisions based on CP period distributions. Examples include stochastic sequential optimization [Liu et al., 2013] and optimization with neural networks trained as decision policies [Dowling and Zhang, 2019]. Recently, an approach combining prediction and decision-making, termed ’decision-focused learning’ or contextual optimization [Sadana et al., 2024], has gained traction. This method trains weighting parameters using decision losses from downstream optimization rather than prediction losses, aligning with the optimization tasks to produce effective decisions [Elmachtoub and Grigas, 2022]. However, it has yet to be applied to the CP shaving problem. Moreover, existing studies often assume a single agent or ’price-taker’ condition, neglecting the impact of various agents’ decisions on CP period predictions. This motivates us to incorporate consumer interactions as a critical aspect of CP shaving design.
@@
Related game formulations. Without specific constraints, the CP shaving problem can be abstracted as agents betting on discrete events, with winners sharing benefits based on their bids. This framework mirrors the classic concept of pari-mutuel betting, in which the odds (or CP periods) fluctuate in response to all participants’ betting strategies [Quandt, 1986, Willis, 1964]. Pari-mutuel betting is commonly applied in sports, where optimal betting strategies—both in size and target—are determined by solving discrete decision problems aimed at maximizing expected returns through the best combination of picks [Kenter, 2016, Rosner, 1975].
@@
The CP shaving problem also shares similarities with congestion games by treating peak demand consumption as a form of resource usage, where increased utilization by multiple agents results in higher costs [Rosenthal, 1973]. Congestion games can often be formulated as potential games, guaranteeing the existence of pure-strategy Nash equilibria (NE) even in weighted settings, provided certain conditions on the utility functions are met [Harks and Klimm, 2016]. The primary difference between congestion games and CP shaving is in their cost structures: in congestion games, costs are determined by the aggregate resource usage, whereas in CP shaving, costs are driven by the peak-period demand—a single, critical resource. This makes the problem more analogous to a bottleneck congestion game, where the objective is to optimize the performance of the worst-performing element [Banner and Orda, 2007]. In such settings, pure-strategy NE has also been shown to exist, although they typically yield a non-optimal price of anarchy (PoA) [Busch and Magdon-Ismail, 2009].
@@
Despite the valuable insights provided by these game-theoretic frameworks, the CP shaving problem exhibits key differences. Specifically, demand shifting can negatively impact consumer comfort and lead to profit losses, the cost structure is dictated by peak-period demand rather than the maximum cost across periods, and the overall system demand must remain unchanged. We address these gaps by integrating these considerations into our model and developing a CP shaving game framework.

### v5

@@ Abstract
Power system operators and electric utility companies often impose a coincident peak demand charge on customers when the aggregate system demand reaches its maximum. This charge incentivizes customers to strategically shift their peak usage away from the system’s collective peak, which helps reduce stress on electricity infrastructure. In this paper, we develop a game-theoretic model to analyze how such strategic behavior affects overall system efficiency. We show that depending on the extent of customers’ demand-shifting capabilities, the resulting coincident peak shaving game can exhibit concavity, quasi-concavity with discontinuities, or non-concavity with discontinuities. In a two-agent, two-period setting, we derive closed-form Nash equilibrium solutions for each scenario and generalize our findings to multi-agent contexts. We prove the stability of the equilibrium points and propose an algorithm for computing equilibrium outcomes under all game configurations. Our results indicate that the peak-shaving outcome at the equilibrium of the game model is comparable to the optimal outcome of the natural centralized model. However, there is a significant loss in efficiency. Under quasi-concave and non-concave conditions, this inefficiency grows with increased customer flexibility and larger disparities in marginal shifting costs; we also examine how the number of agents influences system performance. Finally, numerical simulations with real-world applications validate our theoretical insights.
@@
Coincident peak shaving, Bottleneck game, Equilibrium stability, Demand response, Power system operation, Efficiency loss
@@ I Introduction
Managing peak electricity demand is crucial in power system transmission and distribution due to the need to maintain an instantaneous balance between supply and demand. At the transmission level, total demand must not exceed the maximum generation capacity. At the distribution level, where a tree-branched network of transformer substations steps down voltages, the load at each tier must remain below the capacity of the respective transformers. When demand exceeds the system’s generation or substation capacity, operators may be forced to disconnect portions of the grid to cut off the connecting demand to ensure the secure operation of the remaining system, potentially leading to significant financial losses and adverse impacts on life quality [Kirschen and Strbac,, 2018].
@@
With the advent of smart electric meters capable of recording time-dependent electricity usage and the widespread deployment of distributed energy resources (DER), power system operators and electric utility companies are implementing new pricing schemes to promote demand flexibility and alleviate peak demand stress. One such approach is time-of-use (ToU) tariffs [Salt River Project,, 2025], in which utilities set hourly electricity prices to encourage customer to shift consumption: increasing usage during low-priced periods (typically at night) and reducing it during high-priced periods (typically during the day). However, ToU tariffs mainly focus on recovering the cost of energy supply and shifting daily consumption, and they are less effective at curbing peak demand over longer periods, such as on a monthly basis, because the prices are not set high enough to further incentivize significant peak demand reduction [Trabish,, 2016; Burger et al.,, 2020].
@@
Utility companies may also impose direct peak demand charges, billing customers based on their highest power usage with a prefixed peak price over an extended period—typically a month. These charges not only motivate customers to reduce their peak consumption but also help recover distribution grid investment and maintenance costs, as the delivered grid capacity is determined by a customer’s peak demand [Mullendore,, 2017; Shi et al.,, 2017]. For instance, Consolidated Edison, serving New York City, may charge more than \$40 per kW of peak demand, with the peak being determined by the highest two consecutive 15-minute intervals in any monthly billing cycle. This effectively translates to a cost of approximately \$80 per kWh during peak periods, which is significantly higher than the standard energy supply rate of around 20 to 30 cents per kWh [Consolidated Edison, 2025a, ].
@@
Yet, allocating peak demand charges becomes non-trivial when multiple customers share a capacity bottleneck, such as a constrained substation, because the system’s peak does not directly correspond to each individual customer’s peak [Chen et al., 2020b, ]. Charging a customer for their peak demand may be inequitable if the overall system is not under peak stress, and such charges might not accurately represent the true system peak demand. This issue is particularly critical for utility companies, which are often publicly owned or regulated private monopolies with profit margins capped by their investment costs. Therefore, the revenue from peak demand charges should accurately reflect grid investment expenses based on the actual delivery capacity.
@@
Coincident peak (CP) demand charges bill customers based on their usage during the system’s peak period, offering an alternative that more accurately reflects the system condition and the corresponding investment costs. These charges are added to future electricity bills; for example, Texas’s 4CP mechanism applies charges based on a customer’s average demand during the Electric Reliability Council of Texas (ERCOT) system’s peak hour in each of the four summer months (from June to September), with these charges affecting the following year’s bills. Notably, 4CP charges can represent up to 30% of an organization’s monthly electricity costs [Ögelman,, 2016]. In 2023, the 4CP charge was approximately \$67/kWh [Grid Status,, 2024], with the total coincident peak demand across the ERCOT region reaching 83.6 GW [ERCOT, 2025a, ], translating into an estimated \$5.6 billion payments. According to ERCOT’s annual demand response report, the 4CP mechanism facilitated a 3,500 MW demand reduction during peak periods, resulting in savings of approximately \$234.5 million, or 4.2% of the total 4CP-related costs [Potomac Economics,, 2023].
A significant challenge, however, is that the peak period is identified only in hindsight. While numerous academic and industrial efforts have focused on accurately predicting the system’s peak period [Patrick and Wolak,, 2001; Dowling et al.,, 2018; CPower,, 2016], these studies often overlook the interactions among customers, whose collective behavior determines the system peak. This gap naturally motivates a game-based framework and raises a critical question: Is a game-based framework workable for the CP shaving problem, and how does strategic customer behavior affect system efficiency? Addressing this question is the main focus of this paper.
@@ I-A Summary of main contributions and implications
In this work, we formulate a novel game framework for the CP shaving problem. As a first step toward applying a game-theoretic model to address CP shaving challenges, we consider a simplified two-agent, two-period setting, with extensions to more than two agents discussed in Section VI. The two agents represent clusters of customers or dominant oligopoly customers. The rationale for the two-period setting stems from its ability to represent a peak period and an off-peak period [Anunrojwong et al.,, 2024], as identified from historical market data, including demand shifting patterns and price signals. For example, power system peak demand typically occurs in the early evening, with off-peak periods covering the rest of the day. In the 4CP program in Texas, historical records in 2024 suggest that certain periods, such as around 17:00, are more likely to experience CP events [ERCOT, 2025a, ]. Our model focuses on minimizing system peak demand while deemphasizing off-peak performance, aligning with the primary objective of operators and utilities. This simplified framework offers both insight and tractability for analyzing the fundamental interactions between agents and their impact on system performance. Furthermore, as agents can access the system demand profile from publicly available market information [ERCOT, 2025b, ], they can infer the aggregated response of other agents. We assume that each agent observes the other’s aggregated demand distribution, allowing us to isolate strategic interactions from the complexities of predicting individual valuations. By balancing simplicity with realism, this approach lays a solid foundation for understanding CP shaving strategies and agent behaviors.
@@
With this setting in place, we formulate the CP shaving game model in which agents have a baseline demand at each time period. There is a CP charge price that is announced ahead of time. Each agent has the flexibility to shift some of their demand to earlier or later time periods, and there is a penalty term to account for the cost of this demand-shifting. Each agent’s demand-shifting strategy, influenced by baseline demand, penalty costs, and CP charges, determines their consumption in each period. The collective strategies of all agents shape the total consumption in each time period and, hence, the CP period. Agents with larger baseline demand differences between two periods and lower shifting costs exhibit greater flexibility. Depending on these factors, the model may exhibit concave, quasi-concave with discontinuities, or non-concave with discontinuities characteristics. By contrast, the conventional centralized peak shaving model is formulated as a straightforward concave optimization problem.
@@
We then analytically derive the pure strategy Nash equilibrium (NE) for all CP shaving game structures under the two-agent, two-period setting. Our analysis reveals that the NE is determined by the system’s unbalanced demand and the agents’ shifting capabilities. By treating the system as a switched dynamic system, we prove that the equilibrium of the game model is globally uniformly asymptotically stable, regardless of switching, provided all agents’ demand shifting is within the baseline limit. Furthermore, we demonstrate that a gradient-based algorithm with an update rule serving as a finite difference approximation of the asymptotically stable process can compute the equilibrium point.
@@
We analyze the impact of gaming agents’ strategic behaviors by examining the peak shaving effectiveness and the efficiency loss. Our findings analytically show that the peak-shaving outcome at the equilibrium of the CP shaving game aligns with the optimal outcome of the natural centralized model but incurs significant efficiency losses—except in the concave game, where the outcomes are fully equivalent. For quasiconcave and non-concave games, we prove that the efficiency loss increases with the disparities among agents, as measured by their marginal shifting costs. Additionally, under identical system conditions, we show that greater agent flexibility exacerbates efficiency loss.
@@
We also generalize our findings to a model with more than two agents, and prove the uniqueness of NE under concave and quasiconcave conditions. For non-concave CP shaving games, we show that system demand can still be balanced across two periods and derive the NE at the group level of agents. We also establish the stability of the equilibrium point and validate the effectiveness of the gradient-based algorithm. The NE reveals equivalent peak shaving effectiveness between the game and the centralized model while showing that changes in the number of agents influence the game type and, consequently, the efficiency loss.
For industry practitioners and policymakers, our paper offers the following takeaways:
@@
  • We show that the game model has a stable NE that can be computed efficiently, and applying this model to the CP shaving problem achieves peak shaving effectiveness comparable to that of an efficient centralized model. Thus, operators can adopt the game model without concerns about its impact on CP shaving effectiveness.
@@ I-B Related work
To precisely position our study within the existing literature, we elaborate on our work from three perspectives closely related to our CP shaving problem: (i) Load management, (ii) CP demand charge, and (iii) Related game formulations.
@@
Load management. Load management is a border concept covering peak demand reduction, load shifting, and demand response. Many studies, from the utility’s perspective, formulate optimization problems to design pricing mechanisms - such as tariffs or peak demand charges - aimed at minimizing convex costs or maximizing concave profits [Siano,, 2014]. These pricing design approaches often rely on customers’ price response behavior, typically leading to a two-layer framework (or Stackelberg game model): the upper layer determines the pricing strategy, while the lower layer models demand based on customer response behavior [Chen et al., 2020a, ]. The lower layer may involve constructing utility functions to represent customer preferences [Pennings and Smidts,, 2003; Chen and Xu,, 2024] or applying data-driven methods to learn behavior from historical price and consumption data [Jain et al.,, 2017]. From a demand-side perspective, the key challenge lies in distinguishing between baseline demand and observed demand to better understand customers’ demand reduction decision-making behavior [Agrawal and Yücel,, 2022]. Beyond individual customers, the interactive influence of demand strategies has been widely studied through non-cooperative game formulations and generally achieves peak shaving [Joshi et al.,, 2021]. However, while these load management approaches contribute to peak demand reduction, they primarily focus on maximizing the profit of different market participants rather than directly addressing CP demand reduction in accordance with CP charging mechanisms.
@@
CP demand charge. Academic studies and industrial solutions for CP shaving generally focus on predicting the timing of the system peak and making decisions from the customers’ perspective. Prediction involves estimating the CP period as a probabilistic distribution, often using machine learning tools that leverage input features such as historical demand and weather conditions [Carmona et al.,, 2024; Dowling et al.,, 2018]. Following this, decisions are made based on the predictions. Industrial solutions primarily emphasize short-term prediction, employing auto-regressive methods to iteratively update models and issue warning signals to customers [CPower,, 2016]. Academic solutions typically frame the problem as a scheduling task, making deliberate decisions based on CP period distributions. Examples include stochastic sequential optimization [Liu et al.,, 2013] and optimization with neural networks trained as decision policies [Dowling and Zhang,, 2019]. Recently, an approach combining prediction and decision-making, termed ’decision-focused learning’ or contextual optimization [Sadana et al.,, 2024], has gained traction. This method trains weighting parameters using decision losses from downstream optimization rather than prediction losses, aligning with the optimization tasks to produce effective decisions [Elmachtoub and Grigas,, 2022]. However, it has yet to be applied to the CP shaving problem. Moreover, existing studies often assume a single agent or ’price-taker’ condition, neglecting the impact of various agents’ decisions on CP period predictions. This motivates us to incorporate customer interactions as a critical aspect of CP shaving design.
@@
Related game formulations. Without specific constraints, the CP shaving problem can be abstracted as agents betting on discrete events, with winners sharing benefits based on their bids. This framework mirrors the classic concept of pari-mutuel betting, in which the odds (or CP periods) fluctuate in response to all participants’ betting strategies [Quandt,, 1986; Willis,, 1964]. Pari-mutuel betting is commonly applied in sports, where optimal betting strategies—both in size and target—are determined by solving discrete decision problems aimed at maximizing expected returns through the best combination of picks [Kenter,, 2016; Rosner,, 1975].
@@
The CP shaving problem also shares similarities with congestion games by treating peak demand consumption as a form of resource usage, where increased utilization by multiple agents results in higher costs [Rosenthal,, 1973]. Congestion games can often be formulated as potential games, guaranteeing the existence of pure-strategy Nash equilibria (NE) even in weighted settings, provided certain conditions on the utility functions are met [Harks and Klimm,, 2016]. The primary difference between congestion games and CP shaving is in their cost structures: in congestion games, costs are determined by the aggregate resource usage, whereas in CP shaving, costs are driven by the peak-period demand—a single, critical resource. This makes the problem more analogous to a bottleneck congestion game, where the objective is to optimize the performance of the worst-performing element [Banner and Orda,, 2007]. In such settings, pure-strategy NE has also been shown to exist, though they often result in an inefficient price of anarchy - exceeding one [Busch and Magdon-Ismail,, 2009].
@@
Despite the valuable insights provided by these game-theoretic frameworks, the CP shaving problem exhibits key differences. Specifically, demand shifting can negatively impact customer comfort and lead to profit losses, the cost structure is dictated by peak-period demand rather than the maximum cost across periods, and the overall system demand must remain unchanged. We integrate these considerations into the CP shaving game framework developed and analyzed in this paper.